# On the Smallest Size of Internal Collage Systems

### v1

@@ Abstract
A Straight-Line Program (SLP) for a stirng $T$ is a context-free grammar in Chomsky normal form that derives $T$ only, which can be seen as a compressed form of $T$. Kida et al. introduced collage systems [Theor. Comput. Sci., 2003] to generalize SLPs by adding repetition rules and truncation rules. The smallest size $c(T)$ of collage systems for $T$ has gained attention to see how these generalized rules improve the compression ability of SLPs. Navarro et al. [IEEE Trans. Inf. Theory, 2021] showed that $c(T)\in O(z(T))$ and there is a string family with $c(T)\in\Omega(b(T)\log|T|)$, where $z(T)$ is the number of Lempel-Ziv parsing of $T$ and $b(T)$ is the smallest size of bidirectional schemes for $T$. They also introduced a subclass of collage systems, called internal collage systems, and proved that its smallest size $\hat{c}(T)$ for $T$ is at least $b(T)$. While $c(T)\leq\hat{c}(T)$ is obvious, it is unknown how large $\hat{c}(T)$ is compared to $c(T)$. In this paper, we prove that $\hat{c}(T)=\Theta(c(T))$ by showing that any collage system of size $m$ can be transformed into an internal collage system of size $O(m)$ in $O(m^{2})$ time. Thanks to this result, we can focus on internal collage systems to study the asymptotic behavior of $c(T)$, which helps to suppress excess use of truncation rules. As a direct application, we get $b(T)=O(c(T))$, which answers an open question posed in [Navarro et al., IEEE Trans. Inf. Theory, 2021]. We also give a MAX-SAT formulation to compute $\hat{c}(T)$ for a given $T$.
@@ 1 Introduction
Grammar-based compression is a framework of lossless compression whose outcomes are modeled by grammars like Context-free grammars (CFGs). Grammar-based compression is polular because it is not only powerful to model the practical compressors such as LZ78 [15] and RePair [6], but also suitable to design algorithms working directly on compressed data [8]. In this paper, we focus on CFGs in Chomsky normal form that derive a single string, which are called Straight-Line Programs (SLPs) [3].  [^1] The definition of SLPs varies across the literature. For example, SLPs are not necessarily in Chomsky normal form in [8].
@@
Kida et al. [5] introduced collage systems to generalize SLPs by adding repetition rules and truncation rules. For a repetition rule (a.k.a. run-length rule) $X\rightarrow Y^{r}$ with $r>2$, $X_{i}$ is expanded to the $r$-times repeat of $Y$ and we have $\langle X\rangle=\langle Y\rangle^{r}$. For a truncation rule $X\rightarrow Y[b..e)$ with $1\leq b<e\leq|\langle Y\rangle|+1$, $\langle X\rangle$ is obtained by truncating $\langle Y\rangle$ to the substring $\langle Y\rangle[b..e)$, the substring of $\langle Y\rangle$ starting at $b$ and ending at e1e-1.  [^2] In the original paper [5], the substring truncation is implemented by the prefix truncation and the suffix truncation. Collage systems with no truncation rules are also called Run-length SLPs (RLSLPs) [12] and collage systems with no repetition rules are also called composition systems [2]. The size of a collage system is measured by the number of its nonterminals and let $c(T)$ denote the smallest size of collage systems for a string $T$.
@@
Recently, compressibility measures for highly-repetitive strings have been extensively studied [9, 10]. Although computing $c(T)$ for a given string $T$ is NP-hard [4], it has gained attention as a compressibility measure to see how repetition and/or truncation rules improve the compression ability of SLPs. Navarro et al. [11] showed that $c(T)\in O(z(T))$ and there is a string family with $c(T)\in\Omega(b(T)\log|T|)$, where $z(T)$ is the size of LZ76 parsing [7] of $T$ and $b(T)$ is the smallest size of bidirectional schemes [14] for $T$. They also introduced a subclass of collage systems, called internal collage systems: An internal collage system must have production rules so that, for every nonterminal $X$, there is a way to expand the starting nonterminal to get a sequence that contains $X$ without using truncation rules. This restriction helps to suppress excess use of truncation rules. In particular, a general collage system may have a nonterminal $X$ such that $\langle X\rangle$ is not a substring of $T$ but the substring of $\langle X\rangle$ is truncated and used to represent a part of $T$. This is not the case in internal collage systems where every nonterminal $X$ can be reached from the starting nonterminal without truncation rules, meaning that $\langle X\rangle$ appears in $T$ at least once. Moreover, the restriction enables to define a parsing of $T$ of size linear to the size of an internal collage system and prove its smallest size $\hat{c}(T)$ is in $\Omega(b(T))$ [11].
@@
While $c(T)\leq\hat{c}(T)$ is obvious, it is unknown how large $\hat{c}(T)$ is compared to $c(T)$. In this paper, we prove that $\hat{c}(T)=\Theta(c(T))$ by showing that any collage system of size $m$ can be transformed into an internal collage system of size $O(m)$ in $O(m^{2})$ time. Thanks to this result, we can focus on internal collage systems to study the asymptotic behavior of $c(T)$. As a direct application, we get $b(T)=O(c(T))$, which answers an open question posed in [11].
@@
Kawamoto et al. [4] obtained a hardness result for computing $c(T)$. We just remark that exactly the same proof works to get the following result for $\hat{c}(T)$ because truncation rules are not used to solve the reduced problem, and therefore, there are no difference between $c(T)$ and $\hat{c}(T)$ in the reduction.
Theorem 1.
@@
In this paper, we give a MAX-SAT formulation to compute $\hat{c}(T)$ for a given $T$, extending the previous approaches for SLPs [1] and RLSLPs [4]. Since we now know that $\hat{c}(T)=\Theta(c(T))$, this provides a tool to study the behavior of both $\hat{c}(T)$ and $c(T)$ through computing exact values of $\hat{c}(T)$ for concrete strings (even though it runs in a realistic time only for short strings).
@@ 2.1 Basic notation
A factorization of a string $w$ is a sequence $w_{1},w_{2},\dots,w_{h}$ of substrings of $w$ such that $w=w_{1}w_{2}\cdots w_{h}$. Each substring $w_{i}~(1\leq i\leq h)$ is called a factor of the factorization.
@@ 2.2 Collage Systems and Internal Collage Systems
A collage system [5] of size $m$ has $m$ nonterminals. Each nonterminal $X$ derives a single string $\langle X\rangle$ that is defined by a unique rule to expand $X$ of either one of the following forms, where $Y$ and $Z$ are nonterminals:
@@
  • $X\rightarrow Y[b..e)$ with $1\leq b<e\leq|\langle X_{j}\rangle|+1$ is a truncation rule such that $\langle X\rangle=\langle Y\rangle[b..e)$.
We say that $X$ or its rule refers to symbols in the righthand side of the rule. We assume that nonterminals are sorted such that every nonterminal refers to smaller nonterminals, which is always possible because a collage system must not have a reference loop to represent a finite single string. Also, we assume that there is no useless nonterminal, i.e., all nonterminals are involved in the process of expanding the starting nonterminal. A collage system is called internal if, for every nonterminal $X$, the starting nonterminal can be expanded to get a sequence that contains an occurrence of $X$ without using truncation rules. Let $c(T)$ (resp. $\hat{c}(T)$) denote the smallest size of collage systems (resp. internal collage systems) for a string $T$.
@@
Similarly to [11], we define the grammar tree of a collage system as the tree obtained from the binary parse tree by deleting all descendant nodes under any node that is not the leftmost node with the same label. Since the grammar tree has every nonterminal at most once as the label of an internal node, we sometimes identify an internal node by its label.
Proposition 2.
@@ 3 Collage Systems to Internal Collage Systems
Theorem 3.
@@
Proof.
We call a nonterminal reachable if it appears as a node label of the grammar tree, and otherwise unreachable. Then, a collage system is internal if and only if every nonterminal is reachable. Our algorithm processes unreachable nonterminals $X$ in a top-down manner and modifies the collage system to represent $T$ without using $X$. When we get into unreachable nonterminal $X$, all nonterminals above have been processed to be reachable, and therefore, $X$ is referred to only by truncation rules because referring to $X_{i}$ in a rule other than truncation rules implies that $X$ is reachable. Hence, removing $X$ from the collage system is equivalent to removing the truncation rules that refer to $X$. The process of removing these truncation rules depends on the following cases:
@@
Case 0: Case 0 deals with a base case where $X\rightarrow a$ is an atomic rule. For any nonterminal $Q$ that truncates $X$, it means that $\langle Q\rangle=a$ and it is obvious that the truncation rule can be replaced with $Q\rightarrow a$.
@@
  • When $X\rightarrow Y^{r}$ and the interval $[b..e)$ is included in $(r^{\prime}-1)|\langle Y\rangle|+1..r^{\prime}|\langle Y\rangle|]$ for some $r^{\prime}\in[1..r]$.
@@
In either case, it is easy to replace the truncation rule $Q\rightarrow X[b..e)$ with a rule that truncates a smaller nonterminal $Y$ or $Z$. If the smaller nonterminal is unreachable, the added truncation rule will be processed when we get into the smaller nonterminal.
@@
Case 3: Case 3 deals with rules that truncate unreachable nonterminal $X$ with concatenation or repetition rule spanning exactly two nonterminals. We focus on the case with concatenation rule $X\rightarrow YZ$ as the case with repetition rule can be processed similarly. Let $A$ be the set of all rules that truncate $X$ and their truncated intervals are not included in $1..|\langle Y\rangle|]$ nor $|\langle Y\rangle|+1..|\langle X\rangle|]$. Unlike other cases, we process all truncation rules in $A$ at once.
@@
Time complexity: In every case, we can process a truncation rule in $O(1)$ time each. While the truncation rule is completely removed in Case 0, Case 2 and Case 3, what we do in Case 1 is just switching its reference nonterminal to smaller nonterminals in $O(1)$ time. This switching cost may add up to $O(m)$ until we finally remove it in Case 0, Case 2 or Case 3. Since we remove $O(m)$ truncation rules, the algorithm runs in $O(m^{2})$ time. ∎
It follows from Theorem˜3 that $\hat{c}(T)\leq 9c(T)$. Together with the fact that $c(T)\leq\hat{c}(T)$, we get the following theorem:
Theorem 4.
@@ 4 MAX-SAT formulation to compute $\hat{c}(T)$
In this section, we give a MAX-SAT formulation to compute $\hat{c}(T)$ for a given string $T$. The formulation is based on the ICS-factorization $T=F_{1}\cdots F_{h}$ defined for an internal collage system such that $F_{i}$ is the string derived from the $i$-th leaf of its grammar tree. Recall that $h=m-m_{\mathsf{tr}}-\sigma+1$ holds by Proposition˜2, where $m$ is the size and $m_{\mathsf{tr}}$ is the number of truncation rules of the internal collage system. Thus, minimizing the size $m=h+m_{\mathsf{tr}}+\sigma-1$ is reduced to finding a valid ICT-factorization with the fewest factors and truncation rules.
The following Lemma gives a necessary and sufficient condition for a valid ICT-factorization, which extends Lemma 4.1 of [4] for RLSLPs to incorporate truncation rules.
Lemma 5.
@@
As seen above, every interval in $I$ corresponds to a node of the grammar tree, and hence, it is compatible with tree structures. Finally, we prove condition (iii). For any factor $F_{k}$ longer than $1$, we have set $i_{k}$ and $j_{k}$ so that $F_{k}$ is derived from a nonterminal that refers to smaller nonterminals deriving $F_{i_{k}}\cdots F_{j_{k}}$. Since $G$ does not have a reference loop of nonterminals, there is no reference loop of factors, too. Hence, we can assign integer $D_{k}$ to $F_{k}$ so that $D_{k}$ is larger than integers $\{D_{i_{k}},\dots,D_{j_{k}}\}$ assigned to factors it refers to. Assigning 0 to a factor of length $1$, we see that condition (iii) holds.
@@
$(\Leftarrow)$: Given that every factor $F_{k}$ longer than $1$ is categorized into types to satisfy conditions (i), (ii) and (iii). We can define a reference structure of factors according to condition (i), and it is guaranteed by condition (iii) that there is no reference loop of factors. By condition (ii), $I$ can identify the intervals corresponding to nodes of a tree, which are sufficient to define a reference structure of nonterminals. Although not all nodes are identified, we can fill missing internal nodes by adding concatenation rules to build an internal collage system for the ICS-factorization. ∎
@@
We modify the MAX-SAT formulation in [4] for RLSLPs to encode valid ICS-factorizations of smallest internal collage systems based on Lemma˜5. We say that an ICS-factor longer than $1$ refers to the substring $T[s_{i_{k}}..s_{j_{k+1}})$ defined in Lemma˜5 for each type. Let us call an integer $D_{k}$ satisfying condition (iii) of Lemma˜5 the reference depth of factor $F_{k}$ or the characters covered by $F_{k}$. Note that a type-C factor may refer to a string to the right while type-A and type-B factors always refer to a substring to the left, which adds some complication to our formulation compared to the previous one for RLSLPs. We will encode reference depths to ensure that there is no reference loop.
@@
In the next paragraph, we will define several Boolean variables with multiple parameters in subscripts. For any variable $x$ with some missing subscripts specified by $\circ$”, let $\llbracket x\rrbracket$ denote the set of (possibly tuples of) feasible subscripts to fill the missing part(s). For example, $\llbracket\mathit{ref}^{\mathrm{A}}_{\circ\leftarrow\circ,\circ}\rrbracket=\{(i^{\prime},i,\ell)\mid i\in[1..n-1],\ell\in[2..n-i+1],i^{\prime}\in[1..i-\ell],T[i^{\prime}..i^{\prime}+\ell)=T[i..i+\ell)\}$, and for some fixed $i$ and $\ell$, $\llbracket\mathit{ref}^{\mathrm{A}}_{\circ\leftarrow i,\ell}\rrbracket=\{i^{\prime}\mid(i^{\prime},i,\ell)\in\llbracket\mathit{ref}^{\mathrm{A}}_{\circ\leftarrow\circ,\circ}\rrbracket\}$. We also use $\cdot$” to represent anonymous (arbitrary) subscripts, which are filtered out when used in $\llbracket x\rrbracket$. For example, $\llbracket\mathit{ref}^{\mathrm{A}}_{\cdot\leftarrow\circ,\circ}\rrbracket=\{(i,\ell)\mid(i^{\prime},i,\ell)\in\llbracket\mathit{ref}^{\mathrm{A}}_{\circ\leftarrow\circ,\circ}\rrbracket\}$.
@@
For a given text $T[1..n]$ of length $n$, we define Boolean variables as follows to encode Lemma˜5.

### v2

@@ Abstract
A Straight-Line Program (SLP) for a string $T$ is a context-free grammar in Chomsky normal form that derives $T$ only, which can be seen as a compressed form of $T$. Kida et al. introduced collage systems [Theor. Comput. Sci., 2003] to generalize SLPs by adding repetition rules and truncation rules. The smallest size $c(T)$ of collage systems for $T$ has gained attention to see how these generalized rules improve the compression ability of SLPs. Navarro et al. [IEEE Trans. Inf. Theory, 2021] showed that $c(T)\in O(z(T))$ and there is a string family with $c(T)\in\Omega(b(T)\log|T|)$, where $z(T)$ is the number of phrases in the Lempel-Ziv parsing of $T$ and $b(T)$ is the smallest size of bidirectional schemes for $T$. They also introduced a subclass of collage systems, called internal collage systems, and proved that its smallest size $\hat{c}(T)$ for $T$ is at least $b(T)$. While $c(T)\leq\hat{c}(T)$ is obvious, it is unknown how large $\hat{c}(T)$ is compared to $c(T)$. In this paper, we prove that $\hat{c}(T)=\Theta(c(T))$ by showing that any collage system of size $m$ can be transformed into an internal collage system of size $O(m)$ in $O(m^{2})$ time. Thanks to this result, we can focus on internal collage systems to study the asymptotic behavior of $c(T)$, which helps to suppress excess use of truncation rules. As a direct application, we get $b(T)=O(c(T))$, which answers an open question posed in [Navarro et al., IEEE Trans. Inf. Theory, 2021]. We also give a MAX-SAT formulation to compute $\hat{c}(T)$ for a given $T$.
keywords:
Collage Systems; Dictionary-based compression; Compressibility measures
@@ 1 Introduction
Grammar-based compression is a framework of lossless compression whose outcomes are modeled by grammars like Context-free grammars (CFGs). Grammar-based compression is popular because it is not only powerful to model the practical compressors such as LZ78 [Ziv1978Coi] and RePair [1999LarssonM_OfflinDictionBasedCompr_DCC], but also suitable to design algorithms working directly on compressed data [2012Lohrey_AlgorOnSlpComprStrin]. In this paper, we focus on CFGs in Chomsky normal form that derive a single string, which are called Straight-Line Programs (SLPs) [1995KarpinskiRS_PatterMatchForStrinWith_CPM].  [^1] The definition of SLPs varies across the literature. For example, SLPs are not necessarily in Chomsky normal form in [2012Lohrey_AlgorOnSlpComprStrin].
@@
Kida et al. [2003KidaMSTSA_CollagSystemUnifyFramewFor] introduced collage systems to generalize SLPs by adding repetition rules and truncation rules. For a repetition rule (a.k.a. run-length rule) $X\rightarrow Y^{r}$ with $r>2$, $X_{i}$ is expanded to the $r$-times repeat of $Y$ and we have $\langle X\rangle=\langle Y\rangle^{r}$. For a truncation rule $X\rightarrow Y[b..e)$ with $1\leq b<e\leq|\langle Y\rangle|+1$, $\langle X\rangle$ is obtained by truncating $\langle Y\rangle$ to the substring $\langle Y\rangle[b..e)$, the substring of $\langle Y\rangle$ starting at $b$ and ending at e1e-1.  [^2] In the original paper [2003KidaMSTSA_CollagSystemUnifyFramewFor], the substring truncation is implemented by the prefix truncation and the suffix truncation. Collage systems with no truncation rules are also called Run-length SLPs (RLSLPs) [2016NishimotoIIBT_FullyDynamDataStrucFor_MFCS] and collage systems with no repetition rules are also called composition systems [1996GasieniecKPR_EfficAlgorForLempelZip_SWAT]. The size of a collage system is measured by the number of its nonterminals and let $c(T)$ denote the smallest size of collage systems for a string $T$.
@@
Recently, compressibility measures for highly-repetitive strings have been extensively studied [2021Navarro_IndexHighlRepetStrinCollec_I, 2021Navarro_IndexHighlRepetStrinCollec_II]. Although computing $c(T)$ for a given string $T$ is NP-hard [2024KawamotoIKB_HardnOfSmallRlslpAnd_DCC], it has gained attention as a compressibility measure to see how repetition and/or truncation rules improve the compression ability of SLPs. Navarro et al. [2021NavarroOP_ApproxRatioOfOrderParsin] showed that $c(T)\in O(z(T))$ and there is a string family with $c(T)\in\Omega(b(T)\log|T|)$, where $z(T)$ is the size of LZ76 parsing [1976LempelZ_ComplOfFinitSequen_TIT] of $T$ and $b(T)$ is the smallest size of bidirectional schemes [1982StorerS_DataComprViaTexturSubst] for $T$. They also introduced a subclass of collage systems, called internal collage systems: An internal collage system must have production rules so that, for every nonterminal $X$, the starting nonterminal can be expanded to get a sequence that contains $X$ without using truncation rules. This restriction helps to suppress excess use of truncation rules. In particular, a general collage system may have a nonterminal $X$ such that $\langle X\rangle$ is not a substring of $T$ but the substring of $\langle X\rangle$ is truncated and used to represent a part of $T$. This is not the case in internal collage systems where every nonterminal $X$ can be reached from the starting nonterminal without truncation rules, implying that $\langle X\rangle$ appears in $T$ at least once. Moreover, the restriction enables to define a parsing of $T$ of size linear to the size of an internal collage system and prove its smallest size $\hat{c}(T)$ is in $\Omega(b(T))$ [2021NavarroOP_ApproxRatioOfOrderParsin].
@@
While $c(T)\leq\hat{c}(T)$ is obvious, it is unknown how large $\hat{c}(T)$ is compared to $c(T)$. In this paper, we prove that $\hat{c}(T)=\Theta(c(T))$ by showing that any collage system of size $m$ can be transformed into an internal collage system of size $O(m)$ in $O(m^{2})$ time. Thanks to this result, we can focus on internal collage systems to study the asymptotic behavior of $c(T)$. As a direct application, we get $b(T)=O(c(T))$, which answers an open question posed in [2021NavarroOP_ApproxRatioOfOrderParsin].
@@
Kawamoto et al. [2024KawamotoIKB_HardnOfSmallRlslpAnd_DCC] obtained a hardness result for computing $c(T)$. We just remark that exactly the same proof works to get the following result for $\hat{c}(T)$ because truncation rules are not used to solve the reduced problem, and therefore, there are no difference between $c(T)$ and $\hat{c}(T)$ in the reduction.
Theorem 1.1.
@@
In this paper, we give a MAX-SAT formulation to compute $\hat{c}(T)$ for a given $T$, extending the previous approaches for SLPs [2022BannaiGIKKN_ComputNpHardRepetMeasur_ESA] and RLSLPs [2024KawamotoIKB_HardnOfSmallRlslpAnd_DCC]. Since we now know that $\hat{c}(T)=\Theta(c(T))$, this provides a tool to study the behavior of both $\hat{c}(T)$ and $c(T)$ through computing exact values of $\hat{c}(T)$ for concrete strings (although we may be able to process only short strings in a realistic time).
@@ 2.1 Basic notation
A factorization of a string $w$ is a sequence $w_{1},w_{2},\dots,w_{h}$ of substrings of $w$ such that $w=w_{1}w_{2}\cdots w_{h}$. Each substring $w_{i}\penalty 10000\ (1\leq i\leq h)$ is called a factor of the factorization.
@@ 2.2 Collage Systems and Internal Collage Systems
A collage system [2003KidaMSTSA_CollagSystemUnifyFramewFor] of size $m$ has $m$ nonterminals. Each nonterminal $X$ derives a single string $\langle X\rangle$ that is defined by a unique rule to expand $X$ of either one of the following forms, where $Y$ and $Z$ are nonterminals:
@@
  • $X\rightarrow Y[b..e)$ with $1\leq b<e\leq|\langle Y\rangle|+1$ is a truncation rule such that $\langle X\rangle=\langle Y\rangle[b..e)$.
We say that $X$ or its rule refers to symbols in the righthand side of the rule. We assume that nonterminals are sorted such that every nonterminal refers to smaller nonterminals, which is always possible because a collage system must not have a reference loop to represent a finite single string. Also, we assume that there is no useless nonterminal, i.e., all nonterminals are involved in the process of expanding the starting nonterminal. A collage system is called internal if, for every nonterminal $X$, the starting nonterminal can be expanded to get a sequence that contains $X$ without using truncation rules. Let $c(T)$ (resp. $\hat{c}(T)$) denote the smallest size of collage systems (resp. internal collage systems) for a string $T$.
@@
See Figure˜1 (resp. Figure˜2) for examples of an internal collage system (resp. non-internal collage system) and their binary parse tree.
Figure 1: An illustration of the binary parse tree (right) for the internal collage system having six rules shown left. The internal nodes are depicted by circles, and the leaves by solid boxes. The characters derived from leaves are depicted with dotted boxes. This is an internal collage system because every nonterminal appears as a node label in the binary parse tree.
Figure 2: An illustration of the binary parse tree (right) for the non-internal collage system having six rules shown left. It is not an internal collage system because $X_{4}$ does not appear as a node label in the binary parse tree.
Similarly to [2021NavarroOP_ApproxRatioOfOrderParsin], we define the grammar tree of a collage system as the tree obtained from the binary parse tree by deleting all descendant nodes under any node that is not the leftmost node with the same label (see Figure˜3 for an example). [^3] The idea of grammar trees for internal collage systems is extended from partial parse trees for SLPs [2003Rytter_ApplicOfLempelZivFactor]. Since the grammar tree has every nonterminal at most once as the label of an internal node, we sometimes identify an internal node by its label.
Proposition 2.1.
@@
Figure 3: An illustration of the binary parse tree (right above) and the grammar tree (right below) for the internal collage system having 9 rules shown left. Three internal nodes with label $X_{2}$, $X_{3}$ and $X_{6}$, turn into leaves in the grammar tree because they are not the leftmost nodes with their own label. Observe that Proposition˜2.1 holds, i.e., the number of internal nodes is the size $m=9$ of the collage system, and the number of leaves is $m-m_{\mathsf{tr}}-\sigma+1=9-1-2+1=7$, where $\sigma=2$ is the alphabet size and $m_{\mathsf{tr}}=1$ is the number of truncation rules.
@@ 3 Collage Systems to Internal Collage Systems
Figure 4: Illustration for Case 1. The upper parts show where the truncated substring $\langle Q\rangle$ exists in $\langle X\rangle$ and the lower parts show how the converted collage system represents it without using $X$. The intervals in the truncation rules are abbreviated and shown as $[\cdot)$”. The reference of a nonterminal $Q$ is changed from an unreachable nonterminal $X$ to a smaller nonterminal $Y$ or $Z$. If the new reference is unreachable, it will be processed later.
Figure 5: Illustration for Case 2. At most four new nonterminals including $V$ are enough to represent $\langle Q\rangle$. Note that new rules that truncate $Y$ are introduced, but $Y$ is guaranteed to be reachable via $Q$.
Figure 6: Illustration for Case 3, where the set AA (described in the proof of Theorem˜3.1) is assumed to contain four truncation rules with lefthand sides $Q_{1},Q_{2},Q_{3}$ and $Q_{4}$. For each $Q_{i}\penalty 10000\ (1\leq i\leq 4)$, we introduce at most two new nonterminals, which are guaranteed to be reachable via $Q_{i}$. All such nonterminals other than $S$ and $P$ refer to a reachable nonterminal $S$ or $P$. The new truncation rule $S\rightarrow Y[\cdot)$ (resp. $P\rightarrow Z[\cdot)$) will be processed later if $Y$ (resp. $Z$) is unreachable.
Theorem 3.1.
@@
Proof 3.2.
We call a nonterminal reachable if it appears as a node label of the grammar tree, and otherwise unreachable. Then, a collage system is internal if and only if every nonterminal is reachable. Our algorithm processes unreachable nonterminals $X$ in a top-down manner and modifies the collage system to represent $T$ without using $X$. When we get into an unreachable nonterminal $X$, all nonterminals above have been processed to be reachable, and therefore, $X$ is referred to only by truncation rules because referring to $X_{i}$ in a rule other than truncation rules implies that $X$ is reachable. Hence, removing $X$ from the collage system is equivalent to removing the truncation rules that refer to $X$. The process of removing these truncation rules depends on the following cases:
@@
Case 0: Case 0 is the base case where $X\rightarrow a$ is an atomic rule. For any nonterminal $Q$ that truncates $X$, it means that $\langle Q\rangle=a$ and it is obvious that the truncation rule can be replaced with $Q\rightarrow a$.
@@
  • When $X\rightarrow Y^{r}$ and the interval $[b..e)$ is covered by a single $Y$, i.e., $[b..e)$ is included in $(r^{\prime}-1)|\langle Y\rangle|+1..r^{\prime}|\langle Y\rangle|]$ for some $r^{\prime}\in[1..r]$.
@@
In either case, it is easy to replace the truncation rule $Q\rightarrow X[b..e)$ with a rule that truncates a smaller nonterminal $Y$ or $Z$. If the smaller nonterminal is unreachable, the added truncation rule will be processed later when we get into the nonterminal.
@@
Case 3: Case 3 deals with rules that truncate an unreachable nonterminal $X$ with concatenation or repetition rule spanning exactly two nonterminals. We focus on the case with concatenation rule $X\rightarrow YZ$ as the case with repetition rule can be processed similarly. Let $A$ be the set of all rules that truncate $X$ and their truncated intervals are not included in $1..|\langle Y\rangle|]$ nor $|\langle Y\rangle|+1..|\langle X\rangle|]$. Unlike other cases, we process all truncation rules in $A$ at once.
@@
Time complexity: In every case, we can process a truncation rule in $O(1)$ time each. While the truncation rule is completely removed in Case 0, Case 2 and Case 3, what we do in Case 1 is just switching its reference nonterminal to smaller nonterminals in $O(1)$ time. This switching cost may add up to $O(m)$ until we finally remove it in Case 0, Case 2 or Case 3. Since we remove $O(m)$ truncation rules, the algorithm runs in $O(m^{2})$ time.
It follows from Theorem˜3.1 that $\hat{c}(T)\leq 9c(T)$. Together with the fact that $c(T)\leq\hat{c}(T)$, we get the following theorem:
Theorem 3.3.
@@ 4 MAX-SAT formulation to compute $\hat{c}(T)$
In this section, we give a MAX-SAT formulation to compute $\hat{c}(T)$ for a given string $T$. The formulation is based on the ICS-factorization $T=F_{1}\cdots F_{h}$ defined for an internal collage system such that $F_{i}$ is the string derived from the $i$-th leaf of its grammar tree. Recall that $h=m-m_{\mathsf{tr}}-\sigma+1$ holds by Proposition˜2.1, where $m$ is the size and $m_{\mathsf{tr}}$ is the number of truncation rules of the internal collage system. Thus, minimizing the size $m=h+m_{\mathsf{tr}}+\sigma-1$ is reduced to finding a valid ICS-factorization with the fewest factors and truncation rules.
Example 4.1.
The ICS-factorization of $T=\mathtt{babbbabbaab}$ for the internal collage system of Figure˜3 has seven factors with $F_{1}=\mathtt{ba}$, $F_{2}=\mathtt{b}$, $F_{3}=\mathtt{bb}$, $F_{4}=\mathtt{a}$, $F_{5}=\mathtt{b}$, $F_{6}=\mathtt{ba}$ and $F_{7}=\mathtt{ab}$.
The following Lemma gives a necessary and sufficient condition for a valid ICS-factorization, which extends Lemma 4.1 of [2024KawamotoIKB_HardnOfSmallRlslpAnd_DCC] for RLSLPs to incorporate truncation rules.
Lemma 4.2.
@@
As seen above, every interval in $I$ corresponds to a node of the grammar tree, and hence, it is compatible with tree structures. Also, the interval of a type-B factor corresponds to a leaf node labeled with $Y^{r-1}$, which implies that it is not in $I$. Finally, we prove condition (iv). For any factor $F_{k}$ longer than $1$, we have set $i_{k}$ and $j_{k}$ so that $F_{k}$ is derived from a nonterminal that refers to smaller nonterminals deriving $F_{i_{k}}\cdots F_{j_{k}}$. Since $G$ does not have a reference loop of nonterminals, there is no reference loop of factors, too. Hence, we can assign integer $D_{k}$ to $F_{k}$ so that $D_{k}$ is larger than integers $\{D_{i_{k}},\dots,D_{j_{k}}\}$ assigned to factors it refers to. Assigning 0 to a factor of length $1$, we see that condition (iv) holds.
@@
$(\Leftarrow)$: Given that every factor $F_{k}$ longer than $1$ is categorized into types to satisfy conditions (i), (ii), (iii) and (iv). We can define a reference structure of factors according to condition (i). Referring to a type-B factor alone is avoided by condition (iii), and it is guaranteed by condition (iv) that there is no reference loop of factors. By condition (ii), $I$ can identify the intervals corresponding to nodes of a tree, which are sufficient to define a reference structure of nonterminals. Although not all nodes are identified, we can fill missing internal nodes by adding concatenation rules to build an internal collage system for the ICS-factorization.
@@
We modify the MAX-SAT formulation in [2024KawamotoIKB_HardnOfSmallRlslpAnd_DCC] for RLSLPs to encode valid ICS-factorizations of smallest internal collage systems based on Lemma˜4.2. We say that an ICS-factor longer than $1$ refers to the substring $T[s_{i_{k}}..s_{j_{k+1}})$ defined in Lemma˜4.2 for each type. Let us call an integer $D_{k}$ satisfying condition (iii) of Lemma˜4.2 the reference depth of factor $F_{k}$ or the characters covered by $F_{k}$. Note that a type-C factor may refer to a string to the right while type-A and type-B factors always refer to a substring to the left, which adds some complication to our formulation compared to the previous one for RLSLPs. We will encode reference depths to ensure that there is no reference loop.
@@
In the next paragraph, we will define several Boolean variables with multiple parameters in subscripts. For any variable $x$ with some missing subscripts specified by $\circ$”, let $\llbracket x\rrbracket$ denote the set of (possibly tuples of) feasible subscripts to fill the missing part(s). For example, $\llbracket\mathit{ref}^{\mathrm{A}}_{\circ\leftarrow\circ,\circ}\rrbracket=\{(i^{\prime},i,\ell)\mid i\in[1..n-1],\ell\in[2..n-i+1],i^{\prime}\in[1..i-\ell],T[i^{\prime}..i^{\prime}+\ell)=T[i..i+\ell)\}$, and for some fixed $i$ and $\ell$, $\llbracket\mathit{ref}^{\mathrm{A}}_{\circ\leftarrow i,\ell}\rrbracket=\{i^{\prime}\mid(i^{\prime},i,\ell)\in\llbracket\mathit{ref}^{\mathrm{A}}_{\circ\leftarrow\circ,\circ}\rrbracket\}$. We also use $\bullet$” to represent anonymous (arbitrary) subscripts, which are filtered out when used in $\llbracket x\rrbracket$. For example, $\llbracket\mathit{ref}^{\mathrm{A}}_{\bullet\leftarrow\circ,\circ}\rrbracket=\{(i,\ell)\mid(i^{\prime},i,\ell)\in\llbracket\mathit{ref}^{\mathrm{A}}_{\circ\leftarrow\circ,\circ}\rrbracket\}$.
@@
For a given text $T[1..n]$ of length $n$, we define Boolean variables as follows to encode Lemma˜4.2.